OpenStudy (anonymous):

simplify ^3√8 please show work

OpenStudy (anonymous):

is that the cube root of 8?

OpenStudy (vijay):

6 root 2

OpenStudy (anonymous):

yeah its' infront

OpenStudy (anonymous):

2

OpenStudy (anonymous):

how do you get that

OpenStudy (anonymous):

2^3 = 8

OpenStudy (anonymous):

i still don't understand how you get that

OpenStudy (anonymous):

3√8 \[3\sqrt{2*2*2}\] \[3*2\sqrt{2}\] \[6\sqrt{?}\]

OpenStudy (anonymous):

\[6\sqrt{2}\]